# reference.bib

### v2

@@ F.8 Accuracy of Discriminator
In this section, we assess the reliability of the local discriminator with two metric: accuracy (ACC) and mean squared error (MSE). The discriminator is trained on 4000 samples and tested on 700 samples. To validate the capability of a GPT-4 scorer, we constructe 100 samples with manual labeling and obtain the performance of GPT-4 judger on this dataset. According to Table 9, the ACC of local discriminator is above 0.7 for a 4-label classification problem, with MSE approximately 16% of the total scale (0-3). The scores given by GPT-4 judger are also close to the human labels.
Table 9: Accuracy and MSE for local discriminator and GPT-4 judger. The local discriminator is tested on the validation data scored by GPT-4. GPT-4 Judger is evaluated against ground-truth data with manual label.
@@
ACC 0.710 0.800
MSE 0.488 0.200
F.9 Pseudo Prompt Generation Cost
Algorithm 1 suggests that the generator keeps generating fake prompts until the $(\lambda,\mu,\rho)$-privacy is satisfied. Therefore, it’s crucial to consider the number of sampling times it takes for a generator to produce sufficient qualified pseudo prompts.
$1/\mu$ $5$ $10$ $20$ $30$
Number of Samplings $17.7$ $30.1$ $65.3$ $93.5$
Time per Sampling 1.73 seconds
Table 10 presents the number of generations to perform and the time cost for each generation under different levels of $\mu$. It requires to conduct sampling for an average of $3.23$ times the $1/\mu$. Note that $1/\mu$ can be treated as the minimal number of sampling times to achieve the privacy budget when the prompt contains only one private attribute. The average time per sampling is 1.73 seconds, and thus the overall time consumption for each sample ranges from 30 seconds at $\mu=1/5$ to 161 seconds at $\mu=1/30$.
Appendix G Discussion and Future Work
Extension to diverse task: our framework is experimented on the classification and question-answering task. It’s of significant interest to extend our framework to a broad range of NLP tasks, such as paraphrasing, translation, and summarization. A key challenge in this extension lies in designing an appropriate decomposition methodology for diverse text inputs.
Confusion with historical conversation: we currently consider the one-round conversation setting, while in reality the user may engage in multi-round communications with the server. In such cases, a curious attacker could potentially infer private attributes by detecting patterns in the conversation. It is crucial to develop a confusion strategy along with conducting privacy analysis for multi-round conversations.

### v4

@@ C.5 Accuracy of Discriminator
We assess the reliability of the local discriminator with two metric: accuracy (ACC) and mean squared error (MSE). The discriminator is trained on 4000 samples and tested on 700 samples. To validate the capability of a GPT-4 scorer, we constructe 100 samples with manual labeling and obtain the performance of GPT-4 judger on this dataset. According to Table 10, the ACC of local discriminator is above 0.7 for a 4-label classification problem, with MSE approximately 16% of the total scale (0-3). The scores given by GPT-4 judger are also close to the human labels.
Table 10: Accuracy and MSE for local discriminator and GPT-4 judger.
@@
ACC MSE ACC MSE
Metrics 0.710 0.488 0.800 0.200
   The local discriminator is tested on the validation data scored by GPT-4. GPT-4 Judger is evaluated against ground-truth data with manual label.